# Considering momentum dependence of heavy quarks interaction rate in calculating drag coefficient, and its impact on energy loss of b and c quarks

### v1

@@ V Summary and discussions
In this paper, we present a comprehensive analysis of the drag coefficient within the framework of the Fokker-Planck equation. We account for the momentum dependence of heavy quarks scattering rate when calculating the drag coefficient. This is achieved by modeling the energy loss coefficients as a function of momentum. The results are then compared with a simpler form of drag coefficient which is linked to energy loss through constant coefficients. Our findings indicate that by incorporating linear momentum-dependent terms in the drag calculation, we can more accurately identify the radiative and collisional energy losses for particles with varying momenta, and $R_{AA}$ results demonstrate a better agreement with experimental data. The findings show that as momentum increases, the energy loss also rises, and this enhancement is more significant in radiative energy loss. Additionally, our analysis of $b$ and $c$ quarks reveals that for the $c$ quark, radiative energy loss exceeds collisional energy loss. While, for the $b$ quark, due to its higher mass, collisional energy loss is more prominent than radiative energy loss. Nevertheless, accurately determining free parameter of our model in evolution of $b$ quark is challenging due to the limited amount of experimental data. With more experiments in the future and an expanded dataset for the b quark, we will be able to accurately determine the $\alpha$ and $\beta$ coefficients. This will enable us to quantify the influence of particle’s momentum on scattering rates and, consequently, the associated energy loss.

### v3

@@ V Summary and discussions
We have presented a comprehensive analysis of the drag coefficient within the framework of the Fokker–Planck equation. The momentum dependence of the heavy-quark scattering rate has been incorporated by modeling the energy loss parameters as functions of momentum. The resulting drag coefficient has been compared with that obtained using a simplified model based on constant energy loss terms. Our findings have indicated that considering linear momentum-dependent terms into the drag calculation has allowed for a more accurate identification of both radiative and collisional energy losses for particle distribution in momentum space. The calculated $R_{AA}$ values have shown improved agreement with experimental data. Furthermore, the results have demonstrated that energy loss increases with particle momentum, with the enhancement being more pronounced in radiative energy loss. An analysis of bottom (b) and charm (c) quarks has revealed that radiative energy loss dominates for c quarks, whereas collisional energy loss is more significant for b quarks due to their higher mass. However, accurately determining the free parameters in the evolution of b quarks has remained challenging because of limited experimental data. With more experiments in the future and an expanded dataset for b quark, we will be able to accurately determine the momentum-dependent coefficients. This will allow for a more precise quantification of the influence of heavy particle momentum on their energy loss.
We emphasize that the present work should be regarded as a transparent baseline for identifying momentum-dependent trends in heavy-quark energy loss. Extensions incorporating explicit hadronization models and fully dynamic 3+13+1D hydrodynamic backgrounds are under preparation and will further refine the phenomenology. The current results demonstrate that the momentum-dependent drag already leads to observable modifications in the suppression patterns, motivating more comprehensive future studies.